OpenStudy (anonymous):

add: Write your answer as a fraction in simplest form. \[\frac{ 6 }{ -7 }+\frac{ 7 }{ 5 }\]

OpenStudy (anonymous):

\[\frac{ 6 }{ -7 }+\frac{ 7 }{ 5 }\]

hartnn (hartnn):

cross multiply, hint : \(\huge \frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\)